OpenStudy (anonymous):

Simplify the expression -5+i/2i

OpenStudy (anonymous):

the question should read "write \(\frac{-5+i}{2i}\) in standard form as \(a+bi\) "

OpenStudy (anonymous):

\[\frac{-5+i}{2i}\times \frac{i}{i}\]is a start

OpenStudy (anonymous):

you get \[\frac{-5i-1}{-2}\] or \[\frac{1}{2}+\frac{5}{2}i\]

OpenStudy (anonymous):

-5+i/2i -i -i _______ ?

OpenStudy (anonymous):

@satellite73 I am going to remind you of the help rule not answers